# Incremental Composition of Learned Control Barrier Functions in Unknown Environments

### v1

@@ V Conclusion
We presented an online approach for incremental composition of a non-smooth CBF via locally learned CBFs. As a result, our approach allows for the safe exploration of unknown and static environments. We rely on a class of compactly-supported radial basis functions, called Wendland’s functions, whose end behavior guarantee the valid composition of local CBFs. We validated our approach on a feedback-linearizable planar system and a fixed-velocity Dubins car, and compared our approach to existing works.
Though we observe good performance from using Wendland’s functions, we find that their Lipschitz constants can sometimes be too large to certify validity of local CBFs following the conditions in [1], which are sufficient but not necessary. This suggests that we may benefit from a richer class of basis functions, perhaps compactly-supported neural networks [25], or less-conservative conditions for validity. Another avenue of future work is to generalize our approach to handle non-static and stochastic environments.

### v2

@@ V Conclusion
We presented an online approach for the incremental construction of a non-smooth CBF via local CBFs, allowing for safe exploration in an unknown, static environment, by relying on a class of compactly-supported basis functions, whose end behavior guarantees the valid composition of local CBFs. We validated our approach on a feedback-linearizable planar system and a fixed-velocity Dubins car, and compared to existing work. Though we observe good performance with CS-RBFs, we find their Lipschitz constants to be sometimes too large to certify validity for local CBFs via the conditions in [1], which are sufficient but not necessary. Future work may benefit from more expressive basis functions or less-conservative conditions for validity. Another avenue is generalization to dynamic and stochastic environments.